# Mobility Edge for the Anderson Model on Random Regular Graphs

### 15 Mar 2026 v1

@@ 2.3. Resolvent Criteria for Localization and Delocalization
The following lemma is the primary novelty of this paper. It shows that the conclusions of Theorem 2.1 regarding the resolvent of the Anderson model at the root of Bethe lattice imply the corresponding (de)localization behavior for the matrices $W_{N}$. The proof is given in Section 3.
@@
    If there exists $c>0$ such that $\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c$, then delocalization holds at $E$.
@@ 3.1. Resolvent Estimates
    $\sum_{j=1}^{N}\big|R_{ij}(z)\big|^{2}=\frac{\operatorname{Im}G_{ii}}{\eta}.$ (3.2)
We also require the next lemma, on the uniform integrability of $R_{00}$.
@@
Fix $E\in\mathbb{R}$. The set of random variables $\{(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\}_{\eta\in(0,1)}$ is uniformly integrable.
@@
This is immediate from [5, Lemma 3.3] applied to the moments of order $(s+1)/2$. ∎
@@ 3.2. Preliminary Lemmas
Fix $s\in(0,1/2)\cup[1,\infty)$. There exists $C(s)>1$ such that for all $E\in\mathbb{R}$ and $\eta\geq(\ln N)^{-1}$,
@@
For all $M>0$, there exists $c(M)>0$ such that the following holds. Fix $E\in[-M,M]$ and $\eta\in(0,1)$, and set $I=[E-\eta,E+\eta]$. Then
@@
Fix $d\geq 3$, $t>0$, and $z=E+\mathrm{i}\eta\in\mathbb{C}$ with $\eta>0$. There exists a coupling of $W_{N,t}$ and $H_{d,t}$ such that for all $k\in\mathbb{N}$,
$\lim_{N\rightarrow\infty}G_{kk}(z)=R_{00}(z)$ (3.5)
in probability.
Finally, we state a bound on the second moment of $\operatorname{Im}R_{00}(z)$ that is uniform in $\eta$. The proof is given in Section 7.
Lemma 3.6.
Fix $E\in\mathbb{R}$ and $d\geq 20$. Suppose that there exists $c>0$ such that
$\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c.$
@@ 3.3. Localization
$G_{jl}(z)=\sum_{k=1}^{N}\frac{u_{k}(j)u_{k}(l)}{\lambda_{k}-z},\qquad\operatorname{Im}G_{jj}(z)=\sum_{k=1}^{N}\frac{\eta\,|u_{k}(j)|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}.$ (3.8)
@@
$\sum_{k\in\Lambda_{I}}\big|u_{k}(j)\big|^{2}\leq\sum_{k=1}^{N}\frac{2\eta^{2}\big|u_{k}(j)\big|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}=2\eta\operatorname{Im}G_{jj}(z).$
Multiplying by $N/|\Lambda_{I}|$ gives
@@
Next, recalling (3.5) and using $|G_{11}(z)|\leq 1/\eta$ (from (3.1)), the dominated convergence theorem yields
@@
Combining this with (3.11) and taking $\eta_{0}<c_{1}^{s/2}\delta/2$ completes the proof. ∎
@@
Let $\varepsilon\in(0,1)$ and $p,q>1$ be parameters such that $1/p+1/q=1$. By Hölder’s inequality,
$N=\sum_{j=1}^{N}P_{I}(j)=\frac{1}{N}\sum_{j=1}^{N}(NP_{I}(j))^{\varepsilon}(NP_{I}(j))^{1-\varepsilon}\leq\Bigl(N^{\varepsilon p}Q_{I}(\varepsilon p)\Bigr)^{1/p}\,\Bigl(N^{(1-\varepsilon)q}Q_{I}((1-\varepsilon)q)\Bigr)^{1/q}.$
@@
$\varepsilon=\frac{4}{4-s},\qquad p=2-\frac{s}{2}.$
Then $\varepsilon p=s/2$, $(1-\varepsilon)q=2$, and (p/q)=1s/2(p/q)=1-s/2, yielding
@@ 3.4. Delocalization
Using (3.9), we have
@@
Using the concentration bound (3.3) for $r=2$, we obtain
$\lim_{N\to\infty}\mathbb{P}\left(Q_{I}\leq 4c_{1}^{-2}\cdot\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{2}\Big]+\eta\right)=1.$ (3.14)
From (3.5) and together with the deterministic bound $|G_{11}|\leq\eta^{-1}$ (from (3.1)), the dominated convergence theorem gives
@@ 4.1. Preliminary Reduction
For all symmetric $N\times N$ matrices $M$, $v\in\mathbb{R}^{N}$, $z\in\mathbb{H}$ and $s>0$, define
@@
We frequently abbreviate $F_{s}(z)=F_{s}(z,v,M)$, and we set $\tilde{F}_{s}=F_{s}(z,V_{N},-tA_{N})$, so that with this choice of arguments, $S(z)=G(z)$. We establish the following preliminary concentration estimate in Section 4.4 below, after some preliminary estimates on $F_{s}(z)$ (and a truncated counterpart). The first exponential term comes from Gaussian concentration in $V_{N}$, while the second comes from a coarse Azuma–Freedman bound for the graph randomness in $A_{N}$.
@@
    Further, the same bound holds for the corresponding averages with $\operatorname{Im}S_{ii}(z)$ replaced by $\operatorname{Re}S_{ii}(z)$.
@@
    Fix $s\in(0,1/2)$. Then for every $u>0$ and every $\delta\in(0,\eta^{-1}]$,
    $\mathbb{P}\left(\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\geq u+2\delta^{s}\right)\leq 2\exp\left(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\right)+2\exp\left(-\frac{N\eta^{4}}{Ct^{2}s^{2}}u^{2}\right).$ (4.2)
@@
We begin with the case $s\geq 1$. This follows from setting $u=\eta$ in (4.1). Next, for $s\in(0,1/2)$, we may suppose that $\eta\leq 1$ (by decreasing $\eta$ if necessary). Then the estimate follows from setting $u=\eta/2$ and $\delta=(\eta/4)^{1/s}$ in (4.2). ∎
@@ 4.2. Gaussian Concentration
The following Gaussian concentration inequality is standard (see, e.g., [19, Theorem 5.6]). We say a function $f\colon\mathbb{R}^{N}\rightarrow\mathbb{R}$ is $L$-Lipschitz if $|f(x)-f(y)|\leq L\|x-y\|_{2}$ for all $x,y\in\mathbb{R}^{N}$.
@@
$\mathbb{P}\left(f(X)-\mathbb{E}\big[f(X)\big]\geq t\right)\leq\exp\left(-\frac{t^{2}}{2L^{2}}\right).$
@@
We now derive some gradient bounds that are necessary to apply Lemma 4.2. These bounds are uniform in the matrix $M$, and hence may be applied conditionally on $A_{N}$.
@@
Using (4.3) and the bound $\operatorname{Im}S_{ii}(z)\leq\eta^{-1}$ from (3.1),
@@
By the Ward identity (3.2),
$\sum_{i=1}^{N}|G_{ik}(z)|^{2}=\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{1}{\eta^{2}}.$
@@
For every $s\in(0,1/2)$, the map $x\mapsto x^{s}$ is not Lipschitz near $x=0$. We therefore introduce a truncation at level $\delta>0$ and define, for any $\delta\in(0,\eta^{-1}]$,
@@
Fix $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then
$\|\nabla F_{s,\delta}(z)\|_{2}\leq\frac{s}{\eta^{2}\sqrt{N}}\ \delta^{s-1}.$
@@
Let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\|\varphi_{\delta}^{\prime}\|_{\infty}\leq s\delta^{s-1}$. By the chain rule and (4.3),
$\partial_{k}F_{s,\delta}(z)=\frac{1}{N}\sum_{i=1}^{N}\varphi_{\delta}^{\prime}\big(\operatorname{Im}S_{ii}(z)\big)\partial_{k}\big(\operatorname{Im}S_{ii}(z)\big),\qquad\big|\partial_{k}F_{s,\delta}(z)\big|\leq\frac{1}{N}\sum_{i=1}^{N}\|\varphi_{\delta}^{\prime}\|_{\infty}\big|\partial_{k}\operatorname{Im}S_{ii}(z)\big|.$
@@
$\big|\partial_{k}F_{s,\delta}(z)\big|\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}|S_{ik}(z)|^{2}\leq\frac{s\delta^{s-1}}{N}\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{s\delta^{s-1}}{N\eta^{2}}.$
Summing over $k$ gives
$\|\nabla F_{s,\delta}(z)\|_{2}^{2}\leq N\left(\frac{s}{N\eta^{2}}\delta^{s-1}\right)^{2}=\frac{s^{2}}{N\eta^{4}}\delta^{2s-2}.$
Taking square roots yields the claim. ∎
@@ 4.3. Dependence on the Graph
We write $A(\omega)$ for the adjacency matrix of $A(\omega)$; if a loop occurs, it contributes $2$ to the corresponding diagonal entry, so that every row sum is equal to $d$. Let $\widehat{A}_{N}=A(\omega)$ for $\omega$ chosen uniformly from the set of all pairings of $W$, and let
@@
(see, e.g., [35, equation (5)]), and therefore there exists a constant $p_{d}>0$ such that
@@
We also record the Azuma–Hoeffding martingale concentration estimate that we use. See, e.g., [19, Chapter 2].
@@
    For every $s\in(0,1/2)$ and every $\delta\in(0,\eta^{-1}]$,
    $\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{8s|t|}{N\eta^{2}}\,\delta^{s-1}.$
@@
Set $\Delta=-t(A^{\prime}-A)$. Since $A$ and $A^{\prime}$ differ by one switching of two pairs, the matrix $A^{\prime}-A$ has at most eight nonzero entries, each of magnitude at most $1$. In particular,
$\mathrm{rank}(\Delta)\leq 4,\qquad\|\Delta\|\leq 2|t|.$
@@
$S^{\prime}-S=S\Delta S^{\prime}.$
@@
$\mathrm{rank}(S^{\prime}-S)\leq 4,\qquad\|S^{\prime}-S\|\leq\|S\|\,\|\Delta\|\,\|S^{\prime}\|\leq\frac{2|t|}{\eta^{2}}.$
@@
$\frac{1}{N}\sum_{i=1}^{N}\big|S^{\prime}_{ii}-S_{ii}\big|\leq\frac{1}{N}\|S^{\prime}-S\|_{*}\leq\frac{4}{N}\|S^{\prime}-S\|\leq\frac{8|t|}{N\eta^{2}},$
@@
$\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{s}{N}\eta^{-(s-1)}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{8s|t|}{N\eta^{s+1}}.$
@@
Next, let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\|\varphi_{\delta}^{\prime}\|_{\infty}\leq s\delta^{s-1}$, and therefore
$\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{8s|t|}{N\eta^{2}}\,\delta^{s-1}.$
@@
$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}\big[f(A_{N})\big]\big|\geq u\Big)\leq 2\exp\left(-\frac{u^{2}}{CNdL^{2}}\right).$

### 19 Mar 2026 v2

@@ 2.3. Resolvent Criteria for Localization and Delocalization
The following lemma is the primary novelty of this paper. It shows that the conclusions of Theorem 2.1 regarding the resolvent of the Anderson model at the root of the Bethe lattice imply the corresponding (de)localization behavior for the matrices $W_{N}$. The proof is given in Section 3.
@@
    If there exists $c>0$ such that $\liminf_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c)>c$, then delocalization holds at $E$ for $W_{N}$.
@@ 3.1. Resolvent Estimates
    $\sum_{j=1}^{N}\big|G_{ij}(z)\big|^{2}=\frac{\operatorname{Im}G_{ii}}{\eta}.$ (3.2)
We also require the next lemma, on the uniform integrability of $\operatorname{Im}R_{00}$.
@@
Fix $E\in\mathbb{R}$, $t>0$, and $r\in(0,1)$. The set of random variables $\{(\operatorname{Im}R_{00}(E+\mathrm{i}\eta))^{r}\}_{\eta\in(0,1)}$ is uniformly integrable.
@@
Set $\alpha=(1+r)/2$. By [5, Lemma 3.3], applied with $L=0$ and exponent $\alpha$,
$\sup_{\eta\in(0,1)}\mathbb{E}\!\left[\left(\big(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\big)^{r}\right)^{\alpha/r}\right]\leq\sup_{\eta\in(0,1)}\mathbb{E}\bigl[|R_{00}(E+\mathrm{i}\eta)|^{\alpha}\bigr]<\infty.$
Because $\alpha/r>1$, the claim follows from de la Vallée–Poussin’s criterion. ∎
@@ 3.2. Preliminary Lemmas
Fix $s\in(0,\infty)$, $t>0$, and $d\geq 3$. There exists $C(d,s,t)>1$ such that for all $E\in\mathbb{R}$ and $\eta\geq(\ln N)^{-1}$,
@@
Lemma 3.4.
Fix $d\geq 3$, $t>0$, and $z\in\mathbb{H}$. There exists a coupling of $W_{N,t}$ and $H_{d,t}$ such that
$\lim_{N\rightarrow\infty}G_{11}(z)=R_{00}(z)$ (3.4)
in probability.
@@
For all $t,M>0$ and $d\geq 10$, there exists $c(d,t,M)>0$ such that the following holds. Fix $E\in[-M,M]$ and $\eta\in(0,1)$, and set $I=[E-\eta,E+\eta]$. Then
@@
Finally, we state a bound on the second moment of $\operatorname{Im}R_{00}(z)$ that is uniform in $\eta$. The proof is given in Section 7. The condition $d\geq 10$ could likely be improved, but we refrain from pursuing this to avoid unnecessary complications.
@@
Fix $E\in\mathbb{R}$, $d\geq 10$, and $t>0$. Suppose that there exists $c>0$ such that
$\liminf_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c)>c.$
@@ 3.3. Localization
$\operatorname{Im}G_{jj}(z)=\sum_{k=1}^{N}\frac{\eta\,|u_{k}(j)|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}.$ (3.8)
@@
$\sum_{\lambda_{k}\in\Lambda_{I}}\big|u_{k}(j)\big|^{2}\leq\sum_{k=1}^{N}\frac{2\eta^{2}\big|u_{k}(j)\big|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}=2\eta\operatorname{Im}G_{jj}(z).$
By the first bound in Lemma 3.5, we may restrict to the set where $|\Lambda_{I}|\neq 0$. Multiplying by $N/|\Lambda_{I}|$ gives
@@
Next, recalling (3.4) and using $|G_{11}(z)|\leq 1/\eta$ (from (3.1)), uniform integrability yields
@@
Combining this with (3.11) and taking $\eta_{0}<\min\{c_{2},\,c_{1}^{s/2}\delta/2\}$ completes the proof. ∎
@@
By the first bound in Lemma 3.5, we may restrict to the set where $|\Lambda_{I}|\neq 0$. Let $\varepsilon\in(0,1)$ and $p,q>1$ be parameters such that $1/p+1/q=1$. By Hölder’s inequality,
$N=\sum_{j=1}^{N}P_{I}(j)=\frac{1}{N}\sum_{j=1}^{N}(NP_{I}(j))^{\varepsilon}(NP_{I}(j))^{1-\varepsilon}\leq\Bigl(N^{\varepsilon p}Q_{I}(\varepsilon p)\Bigr)^{1/p}\,\Bigl(N^{(1-\varepsilon)q}Q_{I}((1-\varepsilon)q)\Bigr)^{1/q}.$ (3.13)
@@
$\varepsilon=\frac{s}{4-s},\qquad p=2-\frac{s}{2}.$
Then $\varepsilon p=s/2$, $(1-\varepsilon)q=2$, and (p/q)=1s/2(p/q)=1-s/2. Raising both sides of (3.13) to the $p$th power yields
@@ 3.4. Delocalization
Let $c_{1}(E)>0$ be such that (3.5) yields $\lim_{N\to\infty}\mathbb{P}(|\Lambda_{I}|\geq c_{1}N\eta)=1$. We restrict to the event where $|\Lambda_{I}|\geq c_{1}N\eta$. Using (3.9), we have
@@
Using the concentration bound (3.3) for $s=2$, we obtain
$\lim_{N\to\infty}\mathbb{P}\left(Q_{I}\leq 4c_{1}^{-2}\left(\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{2}\Big]+\eta\right)\right)=1.$ (3.15)
From (3.4), together with the deterministic bound $|G_{11}|\leq\eta^{-1}$ (from (3.1)), uniform integrability gives
@@ 4.1. Preliminary Reduction
For all symmetric $N\times N$ matrices $M$, vectors $v\in\mathbb{R}^{N}$, $z\in\mathbb{H}$ and $s>0$, define
@@
We frequently abbreviate $F_{s}(z)=F_{s}(z,v,M)$, and we write $v_{N}=(g_{1},\dots,g_{N})$ and set $\tilde{F}_{s}=F_{s}(z,v_{N},-tA_{N})$, so that with this choice of arguments, $S(z)=G(z)$. We establish the following preliminary concentration estimate in Section 4.4 below, after some preliminary estimates on $F_{s}(z)$ (and a truncated counterpart). The first exponential term comes from Gaussian concentration in $V_{N}$, while the second comes from a coarse Azuma–Hoeffding bound for the graph randomness in $A_{N}$.
@@
    Further, when $s=1$, the same bound holds for the corresponding averages with $\operatorname{Im}S_{ii}(z)$ replaced by $\operatorname{Re}S_{ii}(z)$.
@@
    Fix $s\in(0,1)$. Then for every $u>0$ and every $\delta\in(0,\eta^{-1}]$,
    $\mathbb{P}\left(\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\geq u+2\delta^{s}\right)\leq 2\exp\left(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\right)+2\exp\left(-\frac{N\eta^{4}}{Ct^{2}s^{2}}\delta^{2-2s}\,u^{2}\right).$ (4.2)
@@
We begin with the case $s\geq 1$. This follows from setting $u=\eta$ in (4.1). Next, for $s\in(0,1)$, if $\eta\geq 1$, then $0\leq\tilde{F}_{s}(z)\leq\eta^{-s}\leq 1$, and therefore
$\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\leq 1\leq\eta$
deterministically, so the claim is trivial. Thus we may assume $\eta\leq 1$. Then the estimate follows from setting $u=\eta/2$ and $\delta=(\eta/4)^{1/s}$ in (4.2). ∎
@@ 4.2. Gaussian Concentration
The following Gaussian concentration inequality is standard (see, e.g., [20, Theorem 5.6]). We say a function $f\colon\mathbb{R}^{N}\rightarrow\mathbb{R}$ is $L$-Lipschitz if $|f(x)-f(y)|\leq L\|x-y\|_{2}$ for all $x,y\in\mathbb{R}^{N}$.
@@
$\mathbb{P}\left(\Big|f(X)-\mathbb{E}\big[f(X)\big]\Big|\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2L^{2}}\right).$
@@
We now derive some Lipschitz bounds that are necessary to apply Lemma 4.2. These bounds are uniform in the matrix $M$, and hence may be applied conditionally on $A_{N}$.
@@
Using (4.3) and the bound $\operatorname{Im}S_{ii}(z)\leq\eta^{-1}$ (analogous to (3.1)),
@@
By the Ward identity (3.2) for $S$,
$\sum_{i=1}^{N}|S_{ik}(z)|^{2}=\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{1}{\eta^{2}}.$
@@
For every $s\in(0,1)$, the map $x\mapsto x^{s}$ is not Lipschitz near $x=0$. We therefore introduce a truncation at level $\delta>0$ and define, for any $\delta\in(0,\eta^{-1}]$,
@@
Fix $s\in(0,1)$ and $\delta\in(0,\eta^{-1}]$. Then, viewed as a function of $v\in\mathbb{R}^{N}$, the map $F_{s,\delta}(z)$ is Lipschitz and
$\operatorname{Lip}(F_{s,\delta}(z))\leq\frac{s}{\eta^{2}\sqrt{N}}\ \delta^{s-1}.$
@@
Let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\varphi_{\delta}$ is Lipschitz on $0,\infty)$, with
$\operatorname{Lip}(\varphi_{\delta})\leq s\delta^{s-1}.$
Fix $k\in\{1,\dots,N\}$ and $h\in\mathbb{R}$. Then
$\big|F_{s,\delta}(v+he_{k};z)-F_{s,\delta}(v;z)\big|$ $\leq\frac{1}{N}\sum_{i=1}^{N}\big|\varphi_{\delta}(\operatorname{Im}S_{ii}(v+he_{k};z))-\varphi_{\delta}(\operatorname{Im}S_{ii}(v;z))\big|$
$\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}\big|\operatorname{Im}S_{ii}(v+he_{k};z)-\operatorname{Im}S_{ii}(v;z)\big|.$
Since $v\mapsto\operatorname{Im}S_{ii}(v;z)$ is $C^{1}$, the fundamental theorem of calculus yields
$\big|\operatorname{Im}S_{ii}(v+he_{k};z)-\operatorname{Im}S_{ii}(v;z)\big|\leq|h|\,\sup_{t\in[0,1]}\big|\partial_{k}(\operatorname{Im}S_{ii})(v+the_{k};z)\big|.$
Hence
$\big|F_{s,\delta}(v+he_{k};z)-F_{s,\delta}(v;z)\big|\leq\frac{s\delta^{s-1}|h|}{N}\sup_{t\in[0,1]}\sum_{i=1}^{N}\big|\partial_{k}(\operatorname{Im}S_{ii})(v+the_{k};z)\big|.$
@@
$\sum_{i=1}^{N}\big|\partial_{k}(\operatorname{Im}S_{ii})(z)\big|\leq\sum_{i=1}^{N}|S_{ik}(z)|^{2}\leq\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{1}{\eta^{2}}.$
Therefore
$\big|F_{s,\delta}(v+he_{k};z)-F_{s,\delta}(v;z)\big|\leq\frac{s\delta^{s-1}}{N\eta^{2}}|h|.$
So the Lipschitz constant in each coordinate direction is at most $\frac{s\delta^{s-1}}{N\eta^{2}}$.
For arbitrary $v,w\in\mathbb{R}^{N}$, varying the coordinates one at a time gives
$|F_{s,\delta}(v;z)-F_{s,\delta}(w;z)|\leq\frac{s\delta^{s-1}}{N\eta^{2}}\,\|v-w\|_{1}\leq\frac{s\delta^{s-1}}{\eta^{2}\sqrt{N}}\,\|v-w\|_{2}.$
This proves the Lipschitz bound. ∎
@@ 4.3. Dependence on the Graph
We write $A(\omega)$ for the adjacency matrix of $\mathcal{G}(\omega)$; if a loop occurs, it contributes $2$ to the corresponding diagonal entry, so that every row sum is equal to $d$. Let $\widehat{A}_{N}=A(\omega)$ for $\omega$ chosen uniformly from the set of all pairings of $W$, and let
@@
(see, e.g., [36, equation (5)]), and therefore there exists a constant $p_{d}>0$ such that
@@
We also record the Azuma–Hoeffding martingale concentration estimate that we use. See, e.g., [20, Chapter 2].
@@
    For every $s\in(0,1)$ and every $\delta\in(0,\eta^{-1}]$,
    $\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{32s|t|}{N\eta^{2}}\,\delta^{s-1}.$
@@
Set $\Delta=-t(A^{\prime}-A)$. Since $A$ and $A^{\prime}$ differ by one switching of two pairs, the matrix $A^{\prime}-A$ is supported on the at most four vertices incident to the switched half-edges. In particular,
$\mathrm{rank}(\Delta)\leq 4.$
Moreover, $A^{\prime}-A$ has at most eight nonzero entries, each of magnitude at most $2$ (since loops contribute $2$ to the diagonal). Hence every row and every column of $A^{\prime}-A$ has $\ell^{1}$-norm at most $8$, so by the Schur test,
$\|\Delta\|\leq 8|t|.$
@@
$S^{\prime}-S=-S\Delta S^{\prime}.$
@@
$\mathrm{rank}(S^{\prime}-S)\leq 4,\qquad\|S^{\prime}-S\|\leq\|S\|\,\|\Delta\|\,\|S^{\prime}\|\leq\frac{8|t|}{\eta^{2}}.$
@@
$\frac{1}{N}\sum_{i=1}^{N}\big|S^{\prime}_{ii}-S_{ii}\big|\leq\frac{1}{N}\|S^{\prime}-S\|_{*}\leq\frac{4}{N}\|S^{\prime}-S\|\leq\frac{32|t|}{N\eta^{2}},$
@@
$\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{s}{N}\eta^{-(s-1)}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{32s|t|}{N\eta^{s+1}}.$
@@
Next, let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Since $s\in(0,1)$, the map $x\mapsto x^{s}$ is concave on $(0,\infty)$, and therefore $\varphi_{\delta}$ is Lipschitz on $0,\infty)$ with
$\operatorname{Lip}(\varphi_{\delta})\leq s\delta^{s-1}.$
Since $\operatorname{Im}S_{ii},\operatorname{Im}S^{\prime}_{ii}\geq 0$, it follows that
$\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|$ $\leq\frac{1}{N}\sum_{i=1}^{N}\big|\varphi_{\delta}(\operatorname{Im}S^{\prime}_{ii})-\varphi_{\delta}(\operatorname{Im}S_{ii})\big|$
$\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|$
$\leq\frac{32s|t|}{N\eta^{2}}\delta^{s-1}.$
@@
$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}\big[f(A_{N})\big]\big|\geq u\Big)\leq 2\exp\left(-\frac{u^{2}}{CNL^{2}}\right).$